# Excluding cosmetic surgeries on hyperbolic 3-manifolds

### v2

@@ 3.3. Improved bound on cosmetic slopes
Theorem 3.19 should be compared to [15, Theorem 7.29]. The statement of that previous result is nearly the same, except that the lower bound on length is $\max\left\{10.1,\,\sqrt{\frac{2\pi}{{\operatorname{sys}}(N)}+58}\right\}$. Thus, in Theorem 3.19, both the constant term and the variable term have become slightly smaller. In particular, when ${\operatorname{sys}}(N)\geq 0.145$, ruling out cosmetic surgeries on $N$ involves checking all slopes shorter than $9.97$ rather than $10.1$. The $1.3\%$ reduction in the normalized length of a slope leads to about a $2.6\%$ reduction in the number of slopes to check. Since the procedure of Section 4 compares each slope shorter than $9.97$ to some number of alternate slopes that may yield a manifold of the same volume, this results in a roughly $4\%$ reduction in the number of pairs of slopes to check. Accordingly, this strengthening is useful in applications.
The proof of Theorem 3.19 closely parallels the proof of [15, Theorem 7.29]. We encourage the reader to go through the arguments of [15, Section 7.3]. While the arguments in that subsection rely on the cone deformation results developed earlier in [15], those results are only used as a black box. Because the argument is so similar, we will mainly point out the few places that have to be modified.
@@
Recall from Definition 3.8 and [15, Remark 4.23] that ${\operatorname{haze}}$ and ${\operatorname{haze}}^{-1}$ are both decreasing functions. It follows that ${\ell_{\textup{new}}}={\ell_{\textup{new}}}(L)$ is a decreasing function of $L$, because ${\operatorname{haze}}\left(1.031-\tfrac{17.85}{L^{2}}\right)$ and $\frac{2\pi}{L({\mathbf{s}})^{2}-14.41}$ are both decreasing functions of $L$. By a derivative computation, the function
@@
From here, the proof is extremely similar to the proof of [15, Lemma 7.26]. As in the proof of that proposition, we substitute ${\ell_{\textup{new}}}(L)=\frac{2\pi}{L^{2}-14.41}$ and apply some algebraic manipulation to show that the desired inequality ${\operatorname{sysmin}}_{\textup{new}}(L)<\frac{2\pi}{L^{2}-56}$ is equivalent to
@@
Compare [15, Equation (7.27)]. We have already checked that the left-hand side of (3.22) is decreasing in $L$; meanwhile, the right-hand side is decreasing by a derivative calculation. As in the proof of [15, Lemma 7.26], we show that as $L\to\infty$, the right-hand side limits to $\frac{56-14.41}{2\pi}=6.619\ldots$. Thus the right-hand side is greater than $6.6$ for all $L$. On the other hand, the left-hand side of (3.22) equals $5.614\ldots$ when $L=12$, hence is lower than $5.7$ for all $L\geq 12$. ∎
@@
This follows exactly the same argument as the proof of [15, Theorem 7.28]. That proof uses a function
@@
  • In order to apply [15, Theorem 7.19] in the proof, one needs to check that $\operatorname{len}(\Sigma)\leq 0.0735$. Since $L\geq L_{0}\geq 9.97$, Theorem 3.14 gives the estimate $\operatorname{len}(\Sigma)\leq 0.07339$, which suffices.
@@
This is a corollary of Theorem 3.23, combined with Mostow rigidity and Gabai’s topological rigidity theorem [16]. See the proof of [15, Theorem 7.29] for full details. ∎
We also prove the following strengthened version of [15, Corollary 1.13]:
@@
Let $N$ be a one-cusped hyperbolic manifold. Suppose that slopes $s,t$ are a cosmetic surgery pair for $N$. If the pair is chirally cosmetic, assume furthermore thatthere does not exist a symmetry $\varphi$ of $N$ with $\varphi(s)=t$. Then, up to reordering $s$ and $t$, the following hold:
@@ 2.     $(2)$
    Either $L(t)<9.97$   or   $v\!\left(1-\frac{14.77}{L(t)^{2}}\right)>\operatorname{vol}(N)-\operatorname{vol}(N(s))$.
Here, $v(z)$ is the decreasing function of Definition 3.9, and we are using the convention that the volume of a non-hyperbolic manifold is 00.
We remark that an orientation–preserving symmetry $\varphi\colon N\to N$ must preserve the homological longitude of $N$ (compare Definition 3.3), hence cannot send a slope $s$ to a distinct slope $t$. This is why the theorem statement is only concerned with symmetries of $N$ in the chirally cosmetic setting.
@@
Suppose without loss of generality that $L(s)\leq L(t)$. Since $s$ and $t$ form a cosmetic surgery pair, and there does not exist a symmetry of $N$ with $\varphi(s)=t$, Theorem 3.19 says the shorter slope $s$ must satisfy (1).
Now, consider slope $t$. If $L(t)<9.97$, then conclusion (2) holds. Alternately, if $L(t)\geq 9.97$, then Theorem 3.14 implies $N(t)$ is hyperbolic, hence $N(s)$ is hyperbolic as well. Theorem 3.14 also implies that
$\operatorname{vol}(N)-\operatorname{vol}(N(s))\>=\>\operatorname{vol}(N)-\operatorname{vol}(N(t))\><\>v\!\left(1-\frac{14.77}{L(t)^{2}}\right),$
@@
Theorem 3.24 says a cosmetic surgery pair on a one-cusped manifold must come from an explicit, finite list. Indeed, there are only finitely many slopes $s$ satisfying (1). There are also finitely many slopes $t$ satisfying $L(t)<9.97$. For each $s$, a theorem of Gromov and Thurston [43, Theorem 6.5.6] says that $\operatorname{vol}(N)-\operatorname{vol}(N(s))>0$. Thus $v\big{(}1-\frac{14.77}{L(t)^{2}}\big{)}$ is bounded below by a strictly positive quantity. Since $v(z)$ decreases toward 00 as $z\to 1$ (compare Definition 3.9), it follows that $\frac{14.77}{L(t)^{2}}$ is bounded below, hence $L(t)$ is bounded above. Thus only finitely many slopes $t$ can satisfy (2).
If $N$ is a multi-cusped hyperbolic manifold, then the statement of Theorem 3.24 still holds with slopes $s,t$ replaced by tuples of slopes ${\mathbf{s}},{\mathbf{s}}^{\prime}$, as in Theorem 3.19. However, there can be infinitely many tuples satisfying the given upper bounds on $L({\mathbf{s}})$ and $L({\mathbf{s}}^{\prime})$.
@@
Let $N_{1}$,$N_{2}$ be one-cusped hyperbolic manifolds. Suppose that $s_{1},s_{2}$ are slopes on $N_{1},N_{2}$ respectively such that $N_{1}(s_{1})$ is homeomorphic to $N_{2}(s_{2})$. Suppose further that there is no homeomorphism of pairs $(N_{1},s_{1})\to(N_{2},s_{2})$. Then the following holds for the set $\{i,j\}=\{1,2\}$:
@@
Here, $v(z)$ is the decreasing function of Definition 3.9, and we are using the convention that the volume of a non-hyperbolic manifold is 00.
@@ 4.1. The procedure
The procedure implemented in our program is not an algorithm, in the sense that the program is not guaranteed to succeed. Nonetheless, any answer produced by the program is trustworthy. All SnapPy and Regina routines used in the program are rigorous, with the exception of length spectra (which are addressed in Step 4 and Remark 4.3).
@@
Next, compute the homological longitude on $\partial C$: that is, a primitive curve ${\lambda_{\rm hom}}$ with $[{\lambda_{\rm hom}}]=0\in H_{1}(N,{\mathbb{Q}})$. As mentioned in Definition 3.3, this curve is unique up to sign.
Next, decide whether $N$ is amphicheiral. This computation can be rigorously certified using a positively oriented ideal triangulation. If $N$ has an orientation-reversing symmetry, then slopes $s$ and $-s$ (in the geometric framing) will always be chirally cosmetic. In this case, we test for purely cosmetic surgeries only.
Finally, if $H_{1}(N,{\mathbb{Z}})\cong{\mathbb{Z}}$, apply Test 2.3 (Casson invariant) to $N$. In particular, if $H_{1}(N,{\mathbb{Z}})\cong{\mathbb{Z}}$ and $\Delta^{\prime\prime}_{N}(1)\neq 0$, then $N$ cannot admit any purely cosmetic surgeries.
@@
By the $6$–theorem of Agol and Lackenby [1, 34], any Dehn filling slope $s$ satisfying $\ell(s)>6$ must produce a hyperbolic manifold $N(s)$. Consequently, all exceptional fillings must come from slopes with $\ell(s)\leq 6$. Using interval arithmetic, SnapPy can compute a set of slopes guaranteed to contain all those of length at most $6$. For each such slope $s$, we attempt to verify that $N(s)$ is hyperbolic by finding a rigorously certified solution to the gluing equations. Failing that, we attempt to rigorously certify that $N(s)$ is exceptional (non-hyperbolic), in one of three ways: by finding obstructions to hyperbolicity in the fundamental group; by finding an incompressible sphere or torus using normal surface theory; or by identifying $N(s)$ as a known Seifert fibered space using Regina.
At the end of this step, the procedure produces a set $\mathcal{E}=\mathcal{E}(N)$ containing certified exceptional slopes and a set $\mathcal{U}=\mathcal{U}(N)$ containing unidentified slopes (whose hyperbolicity could not be settled). The set $\mathcal{U}$ is reported but is not analyzed any further. However, in every instance where we ran our program, we have obtained $\mathcal{U}=\emptyset$. This extends Dunfield’s identification of $\mathcal{E}(N)$ for every $N$ in the SnapPy census [12, Theorem 1.2], with the byproduct that $\mathcal{U}(N)=\emptyset$. Indeed, our code incorporates and extends Dunfield’s code.
@@
For every $s\in\mathcal{E}$, we begin by trying to identify the closed manifold $N(s)$ using the combination of SnapPy and Regina. This means performing the sphere and torus decompositions using Regina, identifying each hyperbolic piece using SnapPy, and identifying each Seifert piece using Regina’s database of closed manifolds. This identification also uses Dunfield’s code [12]. As a result, each non-hyperbolic $N(s)$ is assigned a Regina name that describes the pieces. To save computation time, all of this information is computed just once for each $s\in\mathcal{E}$.
@@
Now, we return to hyperbolic geometry. We begin by computing ${\operatorname{sys}}(N)$. This is the only non-verified hyperbolic geometry invariant that we use in a crucial way. To account for potential numerical error, we subtract $1\%$ from the computed systole. (Compare Remark 4.3.)
@@
For every $p>0$ such that $\mathcal{H}^{p}\neq\emptyset$, and every pair $(s,t)\in\mathcal{H}^{p}\times\mathcal{T}^{p}$ such that $s\neq t$, we try to distinguish $N(s)$ from $N(t)$ using (mostly hyperbolic invariants). We proceed as follows. First, try to distinguish $N(s)$ and $N(t)$ using their volumes, which have been rigorously computed using interval arithmetic. Second, try to distinguish $N(s)$ and $N(t)$ using the Chern–Simons invariant, which can also be computed using interval arithmetic modulo $\pi^{2}/2$. (The Chern–Simons invariant is sensitive to orientation, so this can be particularly useful in the setting where $N(s)$ and $N(t)$ are known to be a chirally cosmetic pair and we are looking for purely cosmetic surgeries.) Third, compare the core homology invariants $\mathcal{D}(N(s),n)$ and $\mathcal{D}(N(t),n)$ for degree $n\leq 7$. If any of the above tests succeeds, we have rigorously distinguished $N(s)$ from $N(t)$.
If the above tests fail, the program tries to distinguish $N(s)$ from $N(t)$ using their complex length spectra up to some cutoff. Unlike the above-mentioned hyperbolic invariants, this is currently nonrigorous, but it still provides suggestive information.
@@
At the end of this process, the program reports any points of failure. First, report the set $\mathcal{U}$ of unidentified slopes (where we could not determine hyperbolicity). Second, report any unequal pairs $(s,t)\in\mathcal{E}\times\mathcal{E}$ that we have failed to distinguish. Third, report any unequal pairs $(s,t)\in\mathcal{H}^{p}\times\mathcal{T}^{p}$ that we have failed to distinguish, and any pairs that have only been distinguished nonrigorously.
Remark 4.3.
All of the steps in the above procedure are rigorous, apart from the computation of systoles and length spectra. The computation of the systole ${\operatorname{sys}}(N)$ for the cusped parent $N$ is a crucial and unavoidable input for all of the subsequent steps. By contrast, the length spectrum of a filled manifold $N(s)$ is used very sparingly, and noted each time it is used.
As described in Step 4, we subtract $1\%$ from the computed value of ${\operatorname{sys}}(N)$ for the cusped parent $N$. While this does not make the computation rigorous, it is extremely conservative. Even when working with standard-precision floats, SnapPy’s calculations are typically correct to more than 14 decimal places.
Our understanding is that a procedure for rigorously verifying the length spectrum of a cusped manifold $N$ is currently in development. Once such a procedure becomes available, the results below will become unconditional.

### v3

@@ 3.3. Improved bound on cosmetic slopes
Theorem 3.19 should be compared to [18, Theorem 7.29]. The statement of that previous result is nearly the same, except that the lower bound on length is $\max\left\{10.1,\,\sqrt{\frac{2\pi}{{\operatorname{sys}}(N)}+58}\right\}$. Thus, in Theorem 3.19, both the constant term and the variable term have become slightly smaller. In particular, when ${\operatorname{sys}}(N)\geq 0.145$, ruling out cosmetic surgeries on $N$ involves checking all slopes shorter than $9.97$ rather than $10.1$. The $\approx 1.3\%$ reduction in the normalized length of a slope leads to about a $\approx 2.6\%$ reduction in the number of slopes to check. Since the procedure of Section 4 compares each slope shorter than $9.97$ to some number of alternate slopes that may yield a manifold of the same volume, this results in a $\approx 4\%$ reduction in the number of pairs of slopes to check. Accordingly, this strengthening is useful in applications.
The proof of Theorem 3.19 closely parallels the proof of [18, Theorem 7.29]. We encourage the reader to go through the arguments of [18, Section 7.3]. While the arguments in that subsection rely on the cone deformation results developed earlier in [18], those results are only used as a black box. Because the argument is so similar, we will mainly point out the few places that have to be modified.
@@
Recall from Definition 3.8 and [18, Remark 4.23] that $\operatorname{haze}$} and ${\operatorname{haze}}^{-1}$ are both decreasing functions. It follows that ${\ell_{\textup{new}}}={\ell_{\textup{new}}}(L)$ is a decreasing function of $L$, because ${\operatorname{haze}}\left(1.031-\tfrac{17.85}{L^{2}}\right)$ and $\frac{2\pi}{L({\mathbf{s}})^{2}-14.41}$ are both decreasing functions of $L$. By a derivative computation, the function
@@
From here, the proof is extremely similar to the proof of [18, Lemma 7.26]. As in the proof of that proposition, we substitute ${\ell_{\textup{new}}}(L)=\frac{2\pi}{L^{2}-14.41}$ and apply some algebraic manipulation to show that the desired inequality ${\operatorname{sysmin}}_{\textup{new}}(L)<\frac{2\pi}{L^{2}-56}$ is equivalent to
@@
Compare [18, Equation (7.27)]. We have already checked that the left-hand side of (3.22) is decreasing in $L$; meanwhile, the right-hand side is decreasing by a derivative calculation. As in the proof of [18, Lemma 7.26], we show that as $L\to\infty$, the right-hand side limits to $\frac{56-14.41}{2\pi}=6.619\ldots$. Thus the right-hand side is greater than $6.6$ for all $L$. On the other hand, the left-hand side of (3.22) equals $5.614\ldots$ when $L=12$, hence is lower than $5.7$ for all $L\geq 12$. ∎
@@
This follows exactly the same argument as the proof of [18, Theorem 7.28]. That proof uses a function
@@
  • In order to apply [18, Theorem 7.19] in the proof, one needs to check that $\operatorname{len}(\Sigma)\leq 0.0735$. Since $L\geq L_{0}\geq 9.97$, Theorem 3.14 gives the estimate $\operatorname{len}(\Sigma)\leq 0.07339$, which suffices.
@@
This is a corollary of Theorem 3.23, combined with Mostow rigidity and Gabai’s topological rigidity theorem [19]. See the proof of [18, Theorem 7.29] for full details. ∎
We also prove the following strengthened version of [18, Corollary 1.13]:
@@
Let $N$ be a one-cusped hyperbolic manifold. Suppose that slopes $s,s^{\prime}$ are a cosmetic surgery pair for $N$. If the pair is chirally cosmetic, assume furthermore that there does not exist a symmetry $\varphi$ of $N$ with $\varphi(s)=s^{\prime}$. Then, up to reordering $s$ and $s^{\prime}$, the following hold:
@@ 2.     $(2)$
    Either $L(s^{\prime})<9.97$   or   $v\!\left(1-\frac{14.77}{L(s^{\prime})^{2}}\right)>\operatorname{vol}(N)-\operatorname{vol}(N(s))$.
Here, $v(z)$ is the decreasing function of Definition 3.9, and we are using the convention that the volume of a non-hyperbolic manifold is 0.
We remark that an orientation–preserving symmetry $\varphi\colon N\to N$ must preserve the homological longitude of $N$ (compare Definition 3.3), hence cannot send a slope $s$ to a distinct slope $s^{\prime}$. This is why the theorem statement is only concerned with symmetries of $N$ in the chirally cosmetic setting.
@@
Suppose without loss of generality that $L(s)\leq L(s^{\prime})$. Since $s$ and $s^{\prime}$ form a cosmetic surgery pair, and there does not exist a symmetry of $N$ with $\varphi(s)=s^{\prime}$, Theorem 3.19 says the shorter slope $s$ must satisfy (1).
Now, consider slope $s^{\prime}$. If $L(s^{\prime})<9.97$, then conclusion (2) holds. Alternately, if $L(s^{\prime})\geq 9.97$, then Theorem 3.14 implies $N(s^{\prime})$ is hyperbolic, hence $N(s)$ is hyperbolic as well. Theorem 3.14 also implies that
$\operatorname{vol}(N)-\operatorname{vol}(N(s))\>=\>\operatorname{vol}(N)-\operatorname{vol}(N(s^{\prime}))\><\>v\!\left(1-\frac{14.77}{L(s^{\prime})^{2}}\right),$
@@
Theorem 3.24 says a cosmetic surgery pair on a one-cusped manifold must come from an explicit, finite list. Indeed, there are only finitely many slopes $s$ satisfying (1). There are also finitely many slopes $s^{\prime}$ satisfying $L(s^{\prime})<9.97$. For each $s$, a theorem of Gromov and Thurston [48, Theorem 6.5.6] says that $\operatorname{vol}(N)-\operatorname{vol}(N(s))>0$. Thus $v\big(1-\frac{14.77}{L(s^{\prime})^{2}}\big)$ is bounded below by a strictly positive quantity. Since $v(z)$ decreases toward 0 as $z\to 1$ (compare Definition 3.9), it follows that $\frac{14.77}{L(s^{\prime})^{2}}$ is bounded below, hence $L(s^{\prime})$ is bounded above. Thus only finitely many slopes $s^{\prime}$ can satisfy (2).
If $N$ is a multi-cusped hyperbolic manifold, then the statement of Theorem 3.24 still holds with slopes $s,s^{\prime}$ replaced by tuples of slopes ${\mathbf{s}},{\mathbf{s}}^{\prime}$, as in Theorem 3.19. However, there can be infinitely many tuples satisfying the given upper bounds on $L({\mathbf{s}})$ and $L({\mathbf{s}}^{\prime})$.
@@
Let $N_{1}$$N_{2}$ be one-cusped hyperbolic manifolds. Suppose that $s_{1},s_{2}$ are slopes on $\partial N_{1},\partial N_{2}$ respectively such that $N_{1}(s_{1})$ is homeomorphic to $N_{2}(s_{2})$. Suppose further that there is no homeomorphism of pairs $(N_{1},s_{1})\to(N_{2},s_{2})$. Then the following holds for the set $\{i,j\}=\{1,2\}$:
@@
Here, $v(z)$ is the decreasing function of Definition 3.9, and we are using the convention that the volume of a non-hyperbolic manifold is 0.
@@ 4.1. The procedure
The procedure implemented in our program is not an algorithm, in the sense that the program is not guaranteed to succeed. Nonetheless, any answer produced by the program is trustworthy. All SnapPy and Regina routines used in the program are rigorous, including rigorously verified intervals for hyperbolic invariants.
@@
Next, compute the homological longitude on $\partial C$: that is, a primitive curve $\lambda_{\rm hom}$} with ${\lambda_{\rm hom}}]=0\in H_{1}(N,{\mathbb{Q}})$. As mentioned in Definition 3.3, this curve is unique up to sign.
Next, decide whether $N$ is amphicheiral. This computation can be rigorously certified using a positively oriented ideal triangulation. If $N$ has an orientation-reversing symmetry, then slopes $s$ and s-s (in the geometric framing) will always be chirally cosmetic. In this case, we test for purely cosmetic surgeries only.
Finally, if $H_{1}(N,{\mathbb{Z}})\cong{\mathbb{Z}}$, apply Test 2.5 (Casson invariant) to $N$. In particular, if $H_{1}(N,{\mathbb{Z}})\cong{\mathbb{Z}}$ and $\Delta^{\prime\prime}_{N}(1)\neq 0$, then $N$ cannot admit any purely cosmetic surgeries.
@@
By the 66–theorem of Agol and Lackenby [1, 39], any Dehn filling slope $s$ satisfying $\ell(s)>6$ must produce a hyperbolic manifold $N(s)$. Consequently, all exceptional fillings must come from slopes with $\ell(s)\leq 6$. Using interval arithmetic, SnapPy can compute a set of slopes guaranteed to contain all those of length at most $6$. For each such slope $s$, we attempt to verify that $N(s)$ is hyperbolic by finding a rigorously certified solution to the gluing equations. Failing that, we attempt to rigorously certify that $N(s)$ is exceptional (non-hyperbolic), in one of three ways: by finding obstructions to hyperbolicity in the fundamental group; by finding an incompressible sphere or torus using normal surface theory; or by identifying $N(s)$ as a known Seifert fibered space using Regina.
At the end of this step, the procedure produces a set $\mathcal{E}=\mathcal{E}(N)$ containing certified exceptional slopes and a set $\mathcal{U}=\mathcal{U}(N)$ containing unidentified slopes (whose hyperbolicity could not be settled). The set $\mathcal{U}$ is reported but is not analyzed any further. However, in every instance where we ran our program, we have obtained $\mathcal{U}=\emptyset$. This extends Dunfield’s identification of $\mathcal{E}(N)$ for every $N$ in the SnapPy census [13, Theorem 1.2], with the byproduct that $\mathcal{U}(N)=\emptyset$. Indeed, our code incorporates and extends Dunfield’s code.
@@
For every $s\in\mathcal{E}$, we begin by trying to identify the closed manifold $N(s)$ using a combination of routines from Regina and SnapPy. This means performing the sphere and torus decompositions using Regina, identifying each hyperbolic piece using SnapPy, and identifying each Seifert piece using Regina’s database of closed manifolds. This identification also uses Dunfield’s code [13]. As a result, each non-hyperbolic $N(s)$ is assigned a Regina name that describes the pieces. To save computation time, all of this information is computed just once for each $s\in\mathcal{E}$.
@@
Now, we return to hyperbolic geometry. We begin by computing ${\operatorname{sys}}(N)$. This can be rigorously certified in SnapPy; see [20].
@@
For every $p>0$ such that $\mathcal{H}^{p}\neq\emptyset$, and every pair $(s,t)\in\mathcal{H}^{p}\times\mathcal{T}^{p}$ such that $s\neq t$, we try to distinguish $N(s)$ from $N(t)$ using (mostly hyperbolic) invariants. We proceed as follows. First, try to distinguish $N(s)$ and $N(t)$ using their volumes, which have been rigorously computed using interval arithmetic. Second, try to distinguish $N(s)$ and $N(t)$ using the Chern–Simons invariant, which can also be computed using interval arithmetic modulo $\pi^{2}/2$. (The Chern–Simons invariant is sensitive to orientation, so this can be particularly useful in the setting where $N(s)$ and $N(t)$ are known to be a chirally cosmetic pair and we are looking for purely cosmetic surgeries.) Third, try to distinguish $N(s)$ from $N(t)$ using their (verified) complex length spectra up to some cutoff. Fourth, compare the core homology invariants $\mathcal{D}(N(s),n)$ and $\mathcal{D}(N(t),n)$ for degree $n\leq 7$. If any of the above tests succeeds, we have rigorously distinguished $N(s)$ from $N(t)$.
@@
At the end of this process, the program reports any points of failure. First, report the set $\mathcal{U}$ of unidentified slopes (where we could not determine hyperbolicity). Second, report any unequal pairs $(s,t)\in\mathcal{E}\times\mathcal{E}$ that we have failed to distinguish. Third, report any unequal pairs $(s,t)\in\mathcal{H}^{p}\times\mathcal{T}^{p}$ that we have failed to distinguish.